# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v1

@@ V Conclusion
  • $\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\leq L_{p}$ (finite sum constraint)
Composite Function Bounds:
  • $\frac{I_{mt}}{I_{mt}+S_{mt}}\in(0,1)$
  • $\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\leq C_{3}$
Lipschitz Constant Derivation:
The partial derivatives of $g(\delta_{mi})$ exist and are bounded on the compact set $0,1]^{L_{p}}$:
$\begin{split}\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}&=\sum_{t=1}^{T}\Bigg[\frac{S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}-I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}}{\big(I_{mt}+S_{mt}\big)^{2}}\cdot\frac{I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\\ &\quad+\frac{I_{mt}}{I_{mt}+S_{mt}}\left(\frac{I_{mt}\,\frac{\partial^{2}S_{mt}}{\partial\delta_{mi}^{2}}-S_{mt}\,\frac{\partial^{2}I_{mt}}{\partial\delta_{mi}^{2}}}{I_{mt}^{2}}-\frac{2\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big(I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big)}{I_{mt}^{3}}\right)\Bigg]\\ &\quad-\chi\Bigg[2\lambda_{1}\,\mathbf{1}_{\{\delta_{mi}-\delta_{mi}^{2}>0\}}\Big((1-2\delta_{mi})^{2}-2(\delta_{mi}-\delta_{mi}^{2})\Big)+2\lambda_{2}\,\mathbf{1}_{\{2-A>0\}}\Bigg].\end{split}$
For all $\delta_{mi}\in[0,1]$, $\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}$ is bounded:
$\left|\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}\right|\leq C_{4}$
Since all component functions are bounded and continuously differentiable on the compact domain, there exists a global Lipschitz constant $L$ such that:
$\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\|\leq L\|\mathbf{x}-\mathbf{y}\|,\quad\forall\mathbf{x},\mathbf{y}\in[0,1]$
This establishes the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, ensuring convergence of the proximal gradient algorithm with appropriate step size selection.
-C Proof of closed-form SINR expression for G-PWPFZF combining:
@@
The desired signal expression in (6) using the combining vectors in (7) and (27).
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Big|\mathbb{E}\Big\{\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mt}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\mathbf{{g}}_{mt}\Big\}\Big|^{2},$ (52)
Using (10) and (28), the (52) can be written as:
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (53)
The first term $\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}$ of interference in (6) can be expanded as:
$\begin{split}\!\!\!\!\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\},\end{split}$ (54)
The first term of (54) is same as (41), the second term of (54) can be written as
$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!(\hat{\mathbf{g}}_{mt}+\tilde{\mathbf{g}}_{mt})\Big|^{2}\!\Bigg\}&{=}\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mt}+\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}\\ &-\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\\ &=\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mt}.\end{split}$ (55)
The expectation in closed-form in third term of (54) can be written as:
$\begin{split}2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Bigg).\end{split}$ (56)
Using (41), (55) and (56) in (54), we get
$\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}=p^{u}_{t}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mt}-\delta_{mi_{t}}\gamma_{mt}).$ (57)
The second term i.e., the pilot contamination term $\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}$ for UE $t$ with pilot sharing UE $k$ in (6), can be evaluated as
$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}\end{split}$ (58)
The first term of (58) is same as (46). The second and third term of (58) can be computed as:
$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}{=}\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mk},$ (59)
$2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg),$ (60)
Using (46), (59) and (60)
@@
$\begin{split}&\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}=p^{u}_{k}\Big|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (61)
The third interference term, for any non-pilot sharing UEs pair $t,k$, in (6) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{UI}_{tk}|^{2}\Big\}&={=}p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &{=}p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(1-\delta_{mi_{k}})\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}(1-\delta_{mi_{k}})\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (62)
The last term of interference in (6) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}&=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (63)
Using the (53), (57), (61), (62) and (63) in (6) gives the SINR closed-form expression in (30).

### v2

@@ V Conclusion and Future Work
  • $\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\leq L_{p}$
  • $\frac{I_{mt}}{I_{mt}+S_{mt}}\in(0,1),\qquad C_{3}\leq\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\leq C_{4}$,
where $C_{1}$, $C_{3}$, $C_{3}$ and $C_{4}$ are finite constants.
The partial derivatives of $g(\delta_{mi})$ is given by
$\begin{split}\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}&=\!\!\sum_{t=1}^{T}\!\!\Bigg[\frac{S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}-I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}}{\big(I_{mt}+S_{mt}\big)^{2}}\frac{I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}+\frac{I_{mt}}{I_{mt}+S_{mt}}\Bigg(\frac{I_{mt}\,\frac{\partial^{2}S_{mt}}{\partial\delta_{mi}^{2}}-S_{mt}\,\frac{\partial^{2}I_{mt}}{\partial\delta_{mi}^{2}}}{I_{mt}^{2}}\\ &-\frac{2\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big(I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big)}{I_{mt}^{3}}\Bigg)\Bigg]-\chi\Bigg[2\lambda_{1}\,\mathbf{1}_{\{\delta_{mi}-\delta_{mi}^{2}>0\}}\Big((1-2\delta_{mi})^{2}\\ &-2(\delta_{mi}-\delta_{mi}^{2})\Big)+2\lambda_{2}\,\mathbf{1}_{\{2-A>0\}}\Bigg].\end{split}$
For all $\delta_{mi}\in[0,1]$, since all component functions $\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}$ are bounded and continuously differentiable on the compact domain $0,1]^{L_{p}}$, the Hessian matrix $\nabla^{2}f(\bm{\delta}_{m})=\nabla g(\bm{\delta}_{m})$ exists and is continuous. Applying the mean value theorem for vector-valued functions, for any $\mathbf{x},\mathbf{y}\in[0,1]^{L_{p}}$, there exists $\mathbf{z}$ on the line segment between $\mathbf{x}$ and $\mathbf{y}$ such that:
$\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\|\leq\|\nabla^{2}f(\mathbf{z})\|_{2}\|\mathbf{x}-\mathbf{y}\|\leq L\|\mathbf{x}-\mathbf{y}\|$
This establishes the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, ensuring convergence of the proximal gradient algorithm.
-C Derivation of closed-form SINR expression for G-PWPFZF combining
@@
The desired signal term in (6)
$|\text{DS}_{t}|^{2}=\Big|\mathbb{E}\Big\{\!\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{{g}}}_{mt}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{{g}}}_{mt}\Big\}\Big|^{2}.$
Using (10) and (28), then
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (C1)
Following the steps in Appendix -A, the first term in (6) can be written as
@@
$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (C2)
Similarly, the last term of interference in (6) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (C3)
Using the (C1), (C2) and (C3) in (6) gives the SINR closed-form expression in (30).